News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 20 October 2000 9.00am EDT

During the last week, Chandra completed the observing schedule as planned.

Last week the Chandra Team were visited by the MSFC Center Director Mr. Art Stephenson. Mr. Stephenson was briefed on the mission status and science highlights, toured the OCC and presented a number of NASA Group Achievement Awards.

A real-time procedure was run to swap to the SSR-B recorder following a loss of approximately 5 hours of data through the ground segment. The data had not yet been overwritten on SSR-A when the loss was discovered and were preserved with the commanded swap to -B. A software patch to modify the SSR-A playback pointer values will allow retrieval of the data, and is in work.

Planned real-time procedures next week include the insertion of the LETG at 6.20pm on 22 Oct and the LETG retraction at 6.20pm EDT on 25 Oct. The team are also preparing for an upcoming lunar eclipse on 27 October during which the solar arrays will be off-pointed to avoid a possible bus over-voltage condition following eclipse exit.

The planned observations for the coming week are as follows.

------------------------------------------
NGC 3227                HRC-S/LETG
AD Leo                  HRC-S/LETG   Oct 24 
KPD0005+5106            HRC-S/LETG   Oct 25 
Radiation Belt
NGC 2541                ACIS-S
0902+343                ACIS-S       Oct 26 
NGC 2683                ACIS-S
3C 212                  ACIS-S
RX J0852.0-4622         ACIS-I
HD100546                ACIS-I
HD104237                ACIS-I
Sgr A                   ACIS-I
GRS 1758-258            HRC-I        Oct 27 
G0.9+0.1                ACIS-S
G208-45                 ACIS-I
NGC 6503                ACIS-S
Lunar Eclipse
Radiation Belt                       Oct 28 
3C263                   ACIS-S
NGC 5457                ACIS-S       Oct 29 
3C294                   ACIS-S
NGC 4414                ACIS-S
NGC 4150                ACIS-S
RXJ0911.4+0551          ACIS-S
IC2560                  ACIS-S

------------------------------------------

All spacecraft subsystems continued to operate nominally.

Return to Status Reports